OpenStudy (anonymous):

what is the minimum value of a+b ; a, b = integers VVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVVV

OpenStudy (anonymous):

what is the minimum value of a+b ; a, b = natural numbers \[\frac{a}{b}=\frac{1}{1^{4}+1^{2}+1}+\frac{2}{2^{4}+2^{2}+1}+...+\frac{100}{100^{4}+100^{2}+1}\]

OpenStudy (anonymous):

@ParthKohli I don't understand. If b = 1, then a must equal the entire sum which is not necessarily closed as a natural number. How can you prove that the sum is closed under natural numbers?

Parth (parthkohli):

Sorry, I didn't see the next sentence.